# Jasmine: Harnessing Diffusion Prior for Self-Supervised Depth Estimation

### v1

@@ 3.3 Scale-Shift GRU
The misalignment of SSI-SI depth.
[…]
@@
where both T and $s_{c}\mathbf{T}$ represent valid relative poses. But if we further introduce a shift $s_{h}$, we can derive the formula to obtain:
@@
where $g_{1}(\cdot)$ and $g_{2}(\cdot)$ are affine invariant transformations. This means that the affine depth of any scene can appear as a plane from a certain perspective (the depth of the plane is $s_{h}$), which is undoubtedly impossible. More detailed derivations are provided in the supplementary material Sec. A.
SD-based MDE predicts Scale-Shift-Invariant Depth $D_{SSI}$. The processed GT depth y mentioned in Sec. 3.1 satisfy the VAE’s inherently boundary [-1, 1], typically obtained by:
$\mathbf{y}=\left(\frac{{D}_{GT}-{D}_{GT,2}}{{D}_{GT,98}-{D}_{GT,2}}-0.5\right)\times 2,$ (11)
where $D_{GT,i}$ corresponds to the $i$percentiles of individual depth maps. Obviously, $\mathbf{y}$ differs from $D_{GT}$ by an absolute scale and shift, which can be recovered by the least squares alignment in evaluation.
@@
Figure 4: Model Structure of Scale Shift GRU. The implementation details are comprehensively described in Sec. 3.3.
Similar to recovering absolute depths, transforming depth distribution requires profound scene understanding, and the scale ($s_{c}$) and shift ($s_{h}$) are tight couplings. So as shown in Fig. 4, the core component Scale-Shift Transformer (SST) employs learnable scale/shift queries that interact with SD’s hidden states (keys/values) via cross-attention. The output vector is split and processed by MLPs to predict $s_{c}$ and $s_{h}$. To enhance spatial awareness, image features $\mathbf{z^{I}}$ are concatenated with depth maps as current input $\mathbf{x^{k}}$. The hidden state $\mathbf{h}^{k}$ evolves through reset $\mathbf{r}^{k+1}$ and update $\mathbf{z}^{k+1}$ gates:
$\begin{array}[]{l}\mathbf{z}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{z}),\mathbf{r}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{r}),\\ \widehat{\mathbf{h}}^{k+1}=\tanh({Conv}([\mathbf{r}^{k+1}\odot\mathbf{h}^{k},\mathbf{x}^{k}],W_{h})),\\ \mathbf{h}^{k+1}=(1-\mathbf{z}^{k+1})\odot\mathbf{h}^{k}+\mathbf{z}^{k+1}\odot\widehat{\mathbf{h}}^{k+1},\end{array}$ (12)
where $k$ denotes iteration steps, and $W_{z},W_{r},W_{h}$ stands for learnable parameters. The refined hidden state $\mathbf{h}^{k+1}$ predicts depth adjustment $D_{\delta}$ to update $D_{k}$, yielding $D_{k+1}$ for subsequent iterations. To balance computational efficiency and GRU’s iterative benefits, we employ two GRU iterations: starting from initial depth $D_{0}$ ($D_{SSI}$) to sequentially produce $D_{1}$ and $D_{2}$ ($D_{SI}$).
GRU is preferred over other architectures due to its reset gate mechanism. This enables the fine-grained $D_{SSI}$ to filter out erroneous supervision signals from reprojection losses. Particularly, in Eq. 12, the reset gate $\mathbf{r}^{k+1}$ identifies regions requiring correction based on current inputs ($\mathbf{h}^{k}$, $\mathbf{x}^{k}$), which are then element-wise multiplied with $\mathbf{h}^{k}$ to compute $\widehat{\mathbf{h}}^{k+1}$. This operation prevents the back-propagation of erroneous gradients to the former step during training. Overall, the SSG not only iteratively aligns SSI depth to SI depth by refining scale-shift parameters but also acts as a gradient filter, which suppresses anomalous gradients caused by artifacts, thereby preserving the fine-grained texture details of SD’s output while enforcing geometric consistency.
@@ 3.4 Steady SD Finetune with Self-Supervision
In the experiment, we observed that the training process was significantly unstable, mainly due to the SD’s enormous size, joint training across modules, and indirect self-supervisory mechanisms. To enhance the convergence reliability and reproducibility, we first introduced a pre-trained self-supervised teacher model (e.g., MonoViT) to estimate $D_{\text{tc}}$ as pseudo labels. $D_{\text{tc}}$ provides direct supervision but has a performance upper bound, which can stabilize model training in the early stages while gradually decreasing loss weights throughout the training process:
where $L_{B}$ is the Berhu Loss, “norm” is the [-1,1] normalization, and “filter” are adaptive strategies to avoid performance bottlenecks, as detailed in the supplementary materials Sec. B.
Furthermore, we implement a module-wise freezing training strategy, which involves 3 key phases: In the beginning, we enable gradient updates for all network components. Subsequently, we freeze the SSG and Posenet modules to decouple depth-pose optimization while maintaining fixed parameters. After achieving convergence to a suboptimal solution, we reintroduce gradient updates to these modules for final optimization. Through extensive experimentation, we demonstrate this module-wise freezing training proves particularly crucial for steady training in complex, multi-module self-supervised systems.
Finally, the total training loss of the Jasmine model is:
@@
where $L_{s}$ refer to Eq.7, $L_{ph}$ refer to Eq. 2, $L_{tc}$ refer to Eq. 13, and $L_{a}$ is some auxiliary adjustment losses (e.g. gds loss[33], edge loss, sky loss) with a tiny weight, and they will be detailed in supplementary material Sec. B.
@@ 4.1 Implement Details
We implement the proposed Jasmine using Accelerate[16] and Pytorch[34] with Stable Diffusion v2[38] as the backbone. Following the pipeline in Fig. 3, we disable text conditioning while maintaining most hyperparameter consistency with E2E FT[32]. The loss weights specified in Sec. 3 are empirically configured as:
$\begin{array}[]{l}\eta_{a}=8\times 10^{-3},\eta_{t1}=0.6,\eta_{t2}=0.9,\eta_{p1}=0.85,\\ \eta_{p2}=0.15,\eta_{\text{step}}=\max(1,30\cdot(\text{step}_{now}/\text{step}_{max})).\end{array}$
Training employs the AdamW optimizer[29] with a base learning rate of $3\times 10^{-5}$. As outlined in Sec. 3.4, the first training phase involved 15k steps, while the second and third phases automatically transitioned, sharing an additional 10k steps for a total of 25k training steps. All experiments are conducted on 8 NVIDIA A800 GPUs with a total batch size of 32, requiring around 1 day. Following [15], we also employed standard data augmentation techniques (horizontal flips, random brightness, contrast, saturation, and hue jitter).
@@ 4.2 Evaluation
Table 1: Quantitative results on the KITTI dataset. For the error-based metrics, the lower value is better; and for the accuracy-based metrics, the higher value is better. The best and second-best results are marked in bold and underline. Jaeho et al* is a combined model that applies Jaeho’s[33] (handle dynamic objects) and TriDepth’s[5] (solve edge flatten) approach. Jasmine* is the same model that uses the least squares alignment. In the data column, Syn, K, and H represent the synthetic, KITTI, and Hypersim datasets, respectively. The number of images and depth labels usage are in brackets. All experiments are conducted at 1024×320 resolution.
@@
E2E FT[32] WACV2025 ZeroShot Syn(74K+74K) 0.112 0.649 4.099 0.180 0.890 0.969 0.985
Lotus[21] ICLR2025 ZeroShot Syn(59K+59K) 0.110 0.611 3.807 0.175 0.892 0.970 0.986
@@
Monodepth2[15] ICCV2019 Mono K(40K+0) 0.115 0.882 4.701 0.190 0.879 0.961 0.982
HR-Depth[30] AAAI2021 Mono K(40K+0) 0.106 0.755 4.472 0.181 0.892 0.966 0.984
R-MSFM6[60] ICCV2021 Mono K(40K+0) 0.108 0.748 4.470 0.185 0.889 0.963 0.982
DevNet[58] ECCV2022 Mono K(40K+0) 0.100 0.699 4.412 0.174 0.893 0.966 0.985
DepthSegNet[31] ECCV2022 Mono K(40K+0) 0.099 0.624 4.165 0.171 0.902 0.969 0.985
SD-SSMDE[36] CVPR2022 Mono K(40K+0) 0.098 0.674 4.187 0.170 0.902 0.968 0.985
MonoViT[56] 3DV 2022 Mono K(40K+0) 0.096 0.714 4.292 0.172 0.908 0.968 0.984
LiteMono[55] CVPR2023 Mono K(40K+0) 0.102 0.746 4.444 0.179 0.896 0.965 0.983
DaCCN[20] ICCV2023 Mono K(40K+0) 0.094 0.624 4.145 0.169 0.909 0.970 0.985
Jaeho et al*.[5, 33] CVPR2024 Mono K(40K+0) 0.091 0.604 4.066 0.164 0.913 0.970 0.986
RPrDepth[19] ECCV2024 Mono K(40K+0) 0.091 0.612 4.098 0.162 0.910 0.971 0.986
Mono-ViFI[28] ECCV2024 Mono K(40K+0) 0.093 0.589 4.072 0.168 0.909 0.969 0.985
@@
4.2.1 Datasets
Unless specified, all datasets are finally resized to 1024$\times$320 resolution for training.
Training Datasets. KITTI[13]: Following the previous work[15], we mainly conduct our experiments on the widely used KITTI dataset. We employ Zhou’s split[59] containing 39,810 training and 4,424 validation samples after removing static frames. The evaluation uses 697 Eigen raw test images with metrics from [15], applying 80m ground truth clipping and Eigen crop preprocessing[8]. Hypersim[37]: This photorealistic synthetic dataset (461 indoor scenes) contributes approximately 28k samples from its official training split for hybrid image reconstruction. Each iteration uses random crops from the original 1024$\times$768 to 1024$\times$320 resolution.
Zero-shot Evaluation Datasets DrivingStereo[49]: Contains 500 images per weather condition (fog, cloudy, rainy, sunny) for zero-shot testing. CityScape[7]: Evaluated on 1,525 test images with dynamic vehicle-rich urban scenes, using ground truth from [47].
HIR Analysis Datasets ETH3D[40]: We resize this high-resolution (6048$\times$4032) dataset to 4K resolution, then randomly cropped to 1024$\times$320 per iteration (898 total samples). Virtual KITTI[4] is a synthetic street scene dataset. We processed this dataset identically to real KITTI data.
4.2.2 Performance Comparison
For all the evaluations, only Jasmine*, E2E FT, and Lotus adopt the least squares alignments. The other self-supervised methods use median alignment. The meanings of each metric are detailed in the supplementary materials.
@@ 4.2.2 Performance Comparison
To fully demonstrate the advantages of our approach, we compare Jasmine against the most efficient SSMDE models and SoTA SD-based methods. As shown in Table 1, Jasmine achieves the best performance across all metrics on the KITTI benchmark. Notably, our method makes significant progress on the $a_{1}$ metric, reflecting an overall improvement in depth estimation accuracy. This systematic advancement stems from the rich prior knowledge of SD. As shown in Fig. 1, without any specialized design for reflective surfaces, our method can accurately distinguish scene elements from their mirror reflections. Additionally, the first row of Fig 5 also demonstrates that our approach preserves structural details better than existing methods. Moreover, when compared to the other SD-based zero-shot models, none of which, including Jasmine*, use the ground truth depth of KITTI, our method demonstrates a substantial performance advantage, further highlighting the value of self-supervised techniques.
@@
Compared to in-domain evaluation, Jasmine exhibits even more remarkable results in zero-shot generalization. Due to the large number of comparison datasets and the unavailability of some models’ open-source code, we only compare with Monodepth2 (the most classic model), MonoViT (renowned for its robustness), MonoViFi (the latest model), and WeatherDepth (trained with additional weather-augmented data, totaling 278k samples), all of which are self-supervised models. In fact, MonoViT already surpasses the zero-shot capability for nearly all models in the SSMDE, making it a sufficiently strong baseline[26]. Additionally, SD-based methods have excelled in generalization, allowing us to conduct a fair zero-shot comparison here.
@@
As shown in Table 2, Jasmine demonstrated state-of-the-art performance on the datasets of Cityscape and four weather scenarios of driving stereo. Remarkably, our method maintains effectiveness in out-of-distribution (OOD) rainy conditions even without specialized training on weather-enhanced datasets like WeatherKITTI[43]. As illustrated in Fig. 1, Jasmine successfully identifies water surface reflections while producing refined depth estimates. The fine-grain estimation details are further demonstrated in Fig. 5, pedestrian chins (second row), tree support structures (third row), and bicycle-rider contours (fourth row) are all predicted delicately and precisely. These sharp results were completely disrupted by the reprojection loss in previous self-supervised methods.
Table 2: Quantitative zero-shot results on the CityScape and DrivingStereo dataset. Alignment protocols and annotation rules follow Table 1’s specifications. The RMSElog and $a_{3}$ metrics are omitted for saving space.
Method AbsRel SqRel RMSE $a_{1}$ $a_{2}$
(a)DrivingStereo
E2E FT[32] 0.160 1.223 5.437 0.795 0.957
Lotus[21] 0.173 1.396 5.816 0.771 0.948
Jasmine* 0.134 0.874 4.666 0.854 0.962
MonoDepth2[15] 0.191 2.865 8.359 0.770 0.923
MonoViT[56] 0.150 1.615 7.657 0.815 0.943
Mono-ViFI[28] 0.158 1.515 6.723 0.798 0.943
WeatherDepth[43] 0.166 1.563 6.986 0.796 0.938
Jasmine 0.136 1.065 5.340 0.850 0.954
(b)Rainy
E2E FT 0.164 1.438 6.671 0.793 0.963
Lotus[21] 0.167 1.398 6.675 0.775 0.961
Jasmine* 0.159 1.487 6.071 0.825 0.957
MonoDepth2[15] 0.260 4.928 12.577 0.609 0.836
MonoViT[56] 0.190 2.574 9.407 0.724 0.913
Mono-ViFI[28] 0.400 7.480 13.960 0.484 0.726
WeatherDepth[43] 0.166 1.874 8.844 0.748 0.942
Jasmine 0.160 1.684 7.194 0.787 0.950
(c)CityScape
E2E FT[32] 0.160 1.512 6.944 0.792 0.955
Lotus[21] 0.147 1.344 6.582 0.824 0.965
Jasmine* 0.107 0.770 5.000 0.907 0.982
MonoDepth2[15] 0.158 1.713 8.185 0.783 0.937
MonoViT[56] 0.140 1.416 7.913 0.802 0.944
Mono-VIFI[28] 0.134 1.290 7.372 0.817 0.950
WeatherDepth[43] 0.137 1.130 6.515 0.837 0.968
Jasmine 0.123 1.037 6.618 0.852 0.967
(d)Cloudy
E2E FT[32] 0.157 1.229 5.522 0.797 0.963
Lotus[21] 0.159 1.263 5.640 0.795 0.959
Jasmine* 0.134 0.910 4.762 0.846 0.968
MonoDepth2[15] 0.192 3.899 10.07 0.775 0.918
MonoViT[56] 0.134 2.028 7.280 0.849 0.955
Mono-VIFI[28] 0.154 1.620 6.883 0.800 0.945
WeatherDepth[43] 0.167 1.956 7.566 0.793 0.934
Jasmine 0.133 1.255 5.651 0.849 0.956
(e)Foggy
E2E FT[32] 0.141 1.089 6.034 0.836 0.977
Lotus[21] 0.150 1.124 6.173 0.798 0.964
Jasmane* 0.113 0.733 4.883 0.897 0.982
MonoDepth2[15] 0.156 2.556 10.425 0.799 0.921
MonoViT[56] 0.107 1.405 7.899 0.882 0.970
Mono-ViFI[28] 0.160 1.715 8.494 0.769 0.932
WeatherDepth[43] 0.132 1.404 7.679 0.859 0.968
Jasmine 0.098 0.876 5.702 0.902 0.977

### v2

@@ 3.3 Scale-Shift GRU
The misalignment of SSI-SI depth. We first analyze the training procedure of SSMDE. Denoting the relative camera poses $\mathcal{T}_{t\rightarrow t^{\prime}}$ as $\textbf{R}|\textbf{T}]$, we can further expand the projproj process in Eq. 1 with:
@@
where both T and $s_{c}\mathbf{T}$ represent valid relative poses. But if we further introduce a shift $s_{h}$, we can derive the formula (detailed derivations at Sec. A) to obtain:
@@
where $g_{1}(\cdot)$ and $g_{2}(\cdot)$ are affine transformations (SSI depth is affine depth). The above equation means that the affine depth of any scene ($g_{1}(D^{\prime})$) can appear as a plane from a certain perspective. This plane has the depth $s_{h}$ and the extrinsic transformation of this perspective and the original one is $\mathbf{R}|\mathbf{T}]$, which is undoubtedly impossible (more explanations in Sec. A). Thus, the shift $s_{h}$ does not exist under the geometric constraints, and SSMDE predicts Scale-Invariant Depth $D_{SI}$.
Additionally, we analyze the training process of SD-based MDE. The processed GT depth y mentioned in Sec. 3.1 satisfies the VAE’svae inherent boundary [-1, 1], typically obtained by:
$\mathbf{y}=\left(({D}_{GT}-{D}_{GT,2})/({D}_{GT,98}-{D}_{GT,2})-0.5\right)\times 2,$ (9)
where $D_{GT,i}$ corresponds to the ii% percentiles of individual depth maps. Obviously, compared to the raw depth, $\mathbf{y}$ is normalized and differs from $D_{GT}$ by an absolute scale and shift, which can be recovered by the least squares alignment in evaluation. Consequently, SD-based MDE predicts Scale-Shift-Invariant Depth $D_{SSI}$.
@@
Transforming the depth distribution from SSI depth to SI depth requires profound scene understanding, and the scale ($s_{c}$) and shift ($s_{h}$) factors are tightly coupled. Therefore, as shown in Fig. 4(a), compared to traditional GRU, SSG introduces a core component Scale-Shift Transformer (SST), and modifies the iterative prediction formula:
$From\quad D_{k+1}=D_{\delta}+D_{k}\quad To\quad D_{k+1}=D_{\delta}+s_{c}\cdot D_{k}+s_{h},$ (10)
where $D_{\delta}=\text{DepthHead}(\mathbf{h}^{k+1})$, $k$ denotes the iteration step and $\mathbf{h}$ denotes the hidden state (preliminaries of GRU in Sec D). Specifically, the SST employs learnable scale/shift queries ($Q_{SC}/Q_{SH}$) that interact with SD’s hidden states (keys/values) via cross-attention. The output vector is subsequently split and processed by MLPs to produce $s_{c}$ and $s_{h}$. For the hidden state update, to enhance spatial awareness, the current input $\mathbf{x}^{k}$ is defined as the concatenation of image features $\mathbf{z}^{I}$ and the current depth $D_{k}$. The hidden state $\mathbf{h}^{k}$ evolves via a standard GRU iteration to produce the refined hidden state $\mathbf{h}^{k+1}$ and subsequently update $D_{k}$ to $D_{k+1}$.
To balance computational efficiency and GRU’s iterative benefits, we employ two GRU iterations: starting from the initial depth $D_{0}$ ($D_{SSI}$) to sequentially produce $D_{1}$ and $D_{2}$ ($D_{SI}$). As shown in Fig. 4(b), the distribution of $D_{0}$ tends to align with the standard SSI depth distribution, while $D_{1}$ and $D_{2}$ progressively converge towards the SI depth. This clearly demonstrates that SSG iteratively aligns SSI depth to SI depth by refining the scale-shift parameters.
Figure 4: (a): Model Structure of SSG. It corresponds to the gray rectangle shown in Fig. 2, standing for an iteration within two consecutive ones. The pipeline of SSG is comprehensively described in Sec. 3.3 (DepthHead is omit in (a) for clear). (b): Depth distribution alignment visualization. We statistically analyze each stage of Jasmine’s SSG module on the KITTI test set. The standard SI and SSI depths are obtained by applying Eq. 9 and dividing the maximum value to the depth GT, respectively.
GRU is preferred over other architectures due to its reset gate mechanism. During training, the reset gate $\mathbf{r}$ can prevent the back-propagation of anomalous gradients to the former step by selectively resetting parts of the hidden state. Therefore, this mechanism enables the fine-grained $D_{SSI}$ to filter out erroneous supervision signals from reprojection losses and exhibit richer details than $D_{SI}$ (shown in Fig. 2). To preserve these fine-grained details in the final prediction, we further constrain the edge alignment between $D_{SSI}$ and $D_{SI}$ with an edge extraction module, detailed in Sec. C.
@@ 3.4 Steady SD Finetune with Self-Supervision
As the first framework to finetune SD with self-supervision, we encountered a novel challenge: training instability, which is mainly due to the SD’s enormous size, joint training across modules, and indirect self-supervisory mechanisms. To enhance the convergence reliability and reproducibility, we explored a straightforward approach by introducing a pre-trained self-supervised teacher model (e.g., MonoViT) to estimate $D_{\text{tc}}$ as pseudo labels. $D_{\text{tc}}$ provides direct supervision but has a performance upper bound, which can stabilize model training in the early stages while gradually decreasing loss weights throughout the training process:
$L_{tc}=\left(L_{B}\left(\text{norm}\left(D_{tc}\right),D_{\text{SSI}}\right)\right.+L_{B}\left(D_{tc},D_{1}\right)\cdot\eta_{t1}\left.+L_{B}\left(\text{filter}\left(D_{tc}\right),D_{SI}\right)\cdot\eta_{t2}\right)/\eta_{\text{step}},$ (11)
where $L_{B}$ is the Berhu Loss, “norm” is the [-1,1] normalization, and “filter” are adaptive strategies to avoid performance bottlenecks. Through extensive experimentation and error bar analysis, we demonstrate that this pseudo-label training proves particularly crucial for steady training in complex, multi-module self-supervised systems. The implementation details are in Sec. C. Finally, the total training loss of the Jasmine model is:
@@
where $L_{s}$ refers to Eq.5, $L_{ph}$ refers to Eq. 2, $L_{tc}$ refers to Eq. 11, and $L_{a}$ is some auxiliary adjustment losses (e.g., gds lossgds $L_{GDS}$, edge loss $L_{e}$, etc.) with a tiny weight. They will be detailed in supplementary material Sec. C.
@@ 4.1 Implement Details
We implement the proposed Jasmine using Accelerateaccelerate and PyTorchpytorch with Stable Diffusion v2sd as the backbone. Following the pipeline in Fig. 2, we disable text conditioning while maintaining most hyperparameter consistency with E2E FTe2eft . The loss weights specified in Sec. 3 are empirically configured as:
$\begin{array}[]{l}\eta_{a}=8e-3,\eta_{t1}=0.6,\eta_{t2}=0.9,\eta_{p1}=0.85,\eta_{p2}=0.15,\eta_{\text{step}}=\max(1,30\cdot(\text{step}_{now}/\text{step}_{max})).\end{array}$
Training uses the AdamW optimizeradamw with a base learning rate of 3e53e-5. All experiments are conducted on 8 NVIDIA A800 GPUs with a total batch size of 32, training for a total of 25k training steps, requiring around 1 day. Following mdp2 , we also employed standard data augmentation techniques (horizontal flips, random brightness, contrast, saturation, and hue jitter).
@@ 4.2 Evaluation
4.2.1 Datasets
[…]
Zero-shot Evaluation Datasets DrivingStereostereo : Contains 500 images per weather condition (fog, cloudy, rainy, sunny) for zero-shot testing. CityScapecity : Evaluated on 1,525 test images with dynamic vehicle-rich urban scenes, using ground truth from manydepth .
[…]
4.2.2 Performance Comparison
Table 1: Quantitative results on the KITTI dataset. For the error-based metrics, the lower value is better; and for the accuracy-based metrics, the higher value is better. The best and second-best results are marked in bold and underline. Jaeho et al* is a combined model that applies both Jaeho’sgds (handle dynamic objects) and TriDepth’stridepth (solve edge flatten) approach. Jasmine* and Jasmine are the same model but use different alignments (discussed in Sec. 4.3). In the data column, Syn, K, and H represent the synthetic, KITTI, and Hypersim datasets, respectively. The number of images and depth labels usage are in brackets. All experiments are conducted at 1024×320 resolution (Performance of Marigold/E2E FT/Lotus is robust to this resolution).
@@ 4.2.2 Performance Comparison
Marigoldmarigold CVPR2024 ZeroShot Syn(74K+74K) 0.120 0.672 4.033 0.184 0.874 0.968 0.985
E2E FTe2eft WACV2025 ZeroShot Syn(74K+74K) 0.112 0.649 4.099 0.180 0.890 0.969 0.985
Lotuslotus ICLR2025 ZeroShot Syn(59K+59K) 0.110 0.611 3.807 0.175 0.892 0.970 0.986
@@
Monodepth2mdp2 ICCV2019 Mono K(40K+0) 0.115 0.882 4.701 0.190 0.879 0.961 0.982
HR-Depthhrdepth AAAI2021 Mono K(40K+0) 0.106 0.755 4.472 0.181 0.892 0.966 0.984
R-MSFM6r-msfm ICCV2021 Mono K(40K+0) 0.108 0.748 4.470 0.185 0.889 0.963 0.982
DevNetdevnet ECCV2022 Mono K(40K+0) 0.100 0.699 4.412 0.174 0.893 0.966 0.985
DepthSegNetdepthsegnet ECCV2022 Mono K(40K+0) 0.099 0.624 4.165 0.171 0.902 0.969 0.985
SD-SSMDEsd-ssmde CVPR2022 Mono K(40K+0) 0.098 0.674 4.187 0.170 0.902 0.968 0.985
MonoViTmonovit 3DV 2022 Mono K(40K+0) 0.096 0.714 4.292 0.172 0.908 0.968 0.984
LiteMonolitemono CVPR2023 Mono K(40K+0) 0.102 0.746 4.444 0.179 0.896 0.965 0.983
DaCCNdaccn ICCV2023 Mono K(40K+0) 0.094 0.624 4.145 0.169 0.909 0.970 0.985
Jaeho et al*.tridepth ; gds CVPR2024 Mono K(40K+0) 0.091 0.604 4.066 0.164 0.913 0.970 0.986
RPrDepthrpr ECCV2024 Mono K(40K+0) 0.091 0.612 4.098 0.162 0.910 0.971 0.986
Mono-ViFIvifi ECCV2024 Mono K(40K+0) 0.093 0.589 4.072 0.168 0.909 0.969 0.985
@@
Table 2: Quantitative zero-shot results on the CityScape and DrivingStereo dataset and its variants (Rainy, Cloudy, Foggy). Alignment protocols and annotation rules follow Table 1’s specifications. AbsRel, RMSE, and $a_{1}$ metrics are shown.
Method DrivingStereo Rainy CityScape Cloudy Foggy
AbsRel RMSE $a_{1}$ AbsRel RMSE $a_{1}$ AbsRel RMSE $a_{1}$ AbsRel RMSE $a_{1}$ AbsRel RMSE $a_{1}$
Marigoldmarigold 0.178 6.638 0.749 0.148 6.770 0.801 0.164 6.632 0.763 0.173 6.881 0.751 0.146 6.545 0.798
E2E FTe2eft 0.160 5.437 0.795 0.164 6.671 0.793 0.160 6.944 0.792 0.157 5.522 0.797 0.141 6.034 0.836
Lotuslotus 0.173 5.816 0.771 0.167 6.675 0.775 0.147 6.582 0.824 0.159 5.640 0.795 0.150 6.173 0.798
Jasmine* 0.134 4.666 0.854 0.159 6.071 0.825 0.107 5.000 0.907 0.134 4.762 0.846 0.113 4.883 0.897
MonoDepth2mdp2 0.191 8.359 0.770 0.260 12.577 0.609 0.158 8.185 0.783 0.192 10.07 0.775 0.156 10.425 0.799
MonoViTmonovit 0.150 7.657 0.815 0.190 9.407 0.724 0.140 7.913 0.802 0.134 7.280 0.849 0.107 7.899 0.882
Mono-ViFIvifi 0.158 6.723 0.798 0.400 13.960 0.484 0.134 7.372 0.817 0.154 6.883 0.800 0.160 8.494 0.769
WeatherDepthweatherdepth 0.166 6.986 0.796 0.166 8.844 0.748 0.137 6.515 0.837 0.167 7.566 0.793 0.132 7.679 0.859
Jasmine 0.136 5.340 0.850 0.160 7.194 0.787 0.123 6.618 0.852 0.133 5.651 0.849 0.098 5.702 0.902
@@
For all the evaluations, only Jasmine*, Marigold, E2E FT, and Lotus adopt the least squares alignment. The other self-supervised methods use median alignment. The definitions and differences of these alignments are detailed discussed in Sec. 4.3. The meaning of each metric is detailed in Sec. G.
@@
To fully demonstrate the advantages of our approach, we compare Jasmine against the most efficient SSMDE models and SoTA SD-based methods. As shown in Table 1, Jasmine achieves the best performance across all metrics on the KITTI benchmark. Notably, our method makes significant progress on the $a_{1}$ metric, reflecting an overall improvement in depth estimation accuracy. This systematic advancement stems from the rich prior knowledge of SD. As shown in Fig. 1, without any specialized design for reflective surfaces, our method can accurately distinguish scene elements from their reflections. Additionally, the first row of Fig. 5 also demonstrates that our approach preserves structural details better than existing methods. Moreover, when compared to the other SD-based zero-shot models, none of which, including Jasmine*, use the ground truth depth of KITTI, our method demonstrates a substantial performance advantage, further highlighting the value of self-supervised techniques.
@@
Compared to in-domain evaluation, Jasmine exhibits even more remarkable results in zero-shot generalization. Due to the large number of comparison datasets and the unavailability of some models’ open-source code, we only compare with Monodepth2 (the most classic model), MonoViT (renowned for its robustness), MonoViFi (the latest model), and WeatherDepth (trained with additional weather-augmented data, totaling 278k samples), all of which are self-supervised models. In fact, MonoViT already surpasses the zero-shot capability for nearly all models in the SSMDE, making it a sufficiently strong baselinerobodepth . Additionally, SD-based methods have excelled in generalization, allowing us to conduct a fair zero-shot comparison here. As shown in Table 2, Jasmine demonstrated state-of-the-art performance on the datasets of Cityscape and four weather scenarios of driving stereo. Remarkably, our method maintains effectiveness in out-of-distribution (OOD) rainy conditions even without specialized training on weather-enhanced datasets like WeatherKITTIweatherdepth . As illustrated in Fig. 1, Jasmine successfully identifies water surface reflections while producing refined depth estimates. The fine-grain estimation details are further demonstrated in Fig. 5, pedestrian chins (second row), tree support structures (third row), and bicycle-rider contours (fourth row) are all predicted delicately and precisely. These sharp results were completely disrupted by the reprojection loss in previous self-supervised methods.
Further Analysis We also deeply compare Jasmine with other SSDE configurations (stereo training and multi-frame inference) and on KITTI improved GT benchmark in Sec. E.2, E.1.
4.3 Analysis of Depth De-normalization
Figure 6: Comparison of different de-normalization schemes. The blue points are predictions after alignment and the green line is the ideal GT depth. Sub-figures (a,c) are the results of LSQ alignment, while (b,d) are median alignment.
The performance gap between Jasmine and Jasmine* in Tables 1 and 2 highlights the impact of de-normalization strategies—median alignment for self-supervised methods and LSQ alignment for zero-shot settings. As the first work bridging these two domains, we provide an in-depth analysis of how these choices affect evaluation.
Depth de-normalization refers to the process of transforming the model’s predicted depth values back to the GT distribution for metric evaluation. The typical de-normalization strategies include:
❶ No operation: Models predict metric depth metric3d : $D_{eval}=D_{pred}$.
❷ Median Alignment: Models predict SI Depth mdp2 , which can be recovered by scaling the ratio of the median of the GT depth to the median of the predicted depth: $D_{eval}=D_{pred}\cdot{\text{median}(D_{GT})}/{\text{median}(D_{pred})}$.
❸ Least-Squares (LSQ) Alignment: Models predict SSI Depth marigold , which can be recovered by affine transformation (scaled and shifted) to best fit the GT depth. The evaluated depth is given by: $D_{eval}=s^{*}D_{pred}+t^{*}$. where the optimal scale $s^{*}$ and shift $t^{*}$ are determined by minimizing the sum of squared differences: $(s^{*},t^{*})=\underset{s,t}{\arg\min}\sum_{i}((s\cdot D_{pred,i}+t)-D_{GT,i})^{2}.$
From Tables 1 and 2, we have following observations:
Metric inconsistency: For the exact same model, the evaluation metrics can vary dramatically depending on the de-normalization method, making direct comparison unfair.
Metrics characteristics: Under LSQ alignment, quadratic metrics (e.g., SqRel, RMSE) are usually better, but first-order metrics (e.g., AbsRel) and overall accuracy ($a_{1}$) are often worse.
Scenario suitability: For in-domain training, median alignment is generally superior, while in zero-shot scenarios, LSQ alignment is usually stronger.
As shown in Fig. 6 (a, c), LSQ alignment tends to accommodate outliers $a$ and $a^{\prime}$, resulting in a larger overall shift in the alignment. In contrast, in sub-figures (b, d), outliers have little effect on the median, so the accuracy for the majority of points is preserved. In the context of depth estimation, $a$ and $a^{\prime}$ can represent the model’s predictions for regions that are difficult to estimate. Clearly, for quadratic metrics, under median alignment, the large error between $a^{\prime}$ and the GT in (d) will be further amplified by squaring, leading to a drop in the overall metric. However, when computing overall accuracy, $a^{\prime}$ or $a$ are typically outside the threshold for $a_{1}$ and thus do not affect it, while the originally accurate predictions for other points become less accurate due to the shift introduced by LSQ alignment. Similarly, first-order metrics also degrade due to these shifts. This explains why Jasmine* achieves better RMSE and SqRel but worse $a_{1}$ and AbsRel compared to Jasmine.
For the last observation, in in-domain scenarios (Table 1), Jasmine outperforms Jasmine* because the shift-free estimation learned by Jasmine is disrupted by LSQ alignment, introducing a suboptimal shift and degrading performance. In out-of-domain scenarios (Table 2), Jasmine* performs better, as the depth distributions of different datasets may differ significantly, and using least-squares to estimate the best fit is clearly a better choice.